solve by method of variation of parameter y'' + y = sec x


1
Expert's answer
2022-02-15T17:32:09-0500

Solution:

The corresponding homogeneous differential equation is



"y''+y=0"

Characteristic (auxiliary) equation



"r^2+1=0""r=\pm i"

The general solution of the homogeneous differential equation is



"y_h=c_1 \cos x+c_2\sin x""y'=c_1 '\cos x-c_1 \sin x+c_2'\sin x+c_2\cos x"

If


"c_1 '\cos x+c_2'\sin x=0,"

then



"y''=-c_1'\sin x-c_1\cos x+c_2'\cos x-c_2\sin x"

Substitute



"-c_1'\sin x-c_1\cos x+c_2'\cos x-c_2\sin x""+c_1 \cos x+c_2\sin x=\dfrac{1}{\cos x}"

We have



"c_1 '\cos x+c_2'\sin x=0,""-c_1'\sin x+c_2'\cos x=\dfrac{1}{\cos x}"





"c_1'=-\dfrac{\sin x}{\cos x}c_2'""\dfrac{\sin^2 x}{\cos x}c_2'+c_2'\cos x=\dfrac{1}{\cos x}"





"c_1'=-\dfrac{\sin x}{\cos x}""c_2'=1"

Integrate



"c_1=-\int \dfrac{\sin x}{\cos x}dx=\ln (|\cos x|)+C_1""c_2=\int dx=x+C_2"

The general solution of the non-homogeneous differential equation is



"y_h=\ln (|\cos x|) \cos x+x\sin x+C_1\cos x+C_2\sin x"